A point charge π‘ž = 8.00 π‘₯ 10βˆ’9𝐢 is at the center of the cube with sides of length 0.200 m. What is the electric flux through one of the six faces of the cube?

1
Expert's answer
2022-02-21T12:54:56-0500

Answer



"\phi=\frac{Q}{\epsilon_0}\\=\frac{8.00 *10^{-9}}{8.85*10^{-12}}\\=903.95Wb"

the electric flux through one of the six faces of the cube

"=\frac{903.95Wb}{6}\\=150.66Wb"